An electrician charges a base fee of $70 plus $50 for each hour of work. Create a table that shows the amount the electrician charges for 1,2,3, and 4 hours of work. Let x represent the number of hours and y represent the amount charged for x hours. Is this relation a function?


1
Expert's answer
2022-01-09T15:59:09-0500

the table is shown



the relation is a function such that

y=x*50......(1)

total= y+70....... (2)

also

total =(x*50)+70 where x number of hours